# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ 5 Conclusion
In this paper, we conduct a systematic evaluation of various defense methods both theoretically and empirically. To perform the evaluation, we introduce a stronger attack method that is robust to a wide range of defense methods in the setting of honest-but-curious server. The method enhances the conventional gradient inversion attack by combining it with the feature reconstruction attack, where a tensor-based attack method is adopted. Our theoretical and empirical findings indicate that gradient pruning is the most effective defense method among many others, including gradient clipping, dropout, adding noise, local aggregation, etc.
Appendix A Preliminaries for Theoretical Analysis
A.1 Tensor Method
@@
Various of noisy tensor decomposition can be used in tensor based attack (Wang et al., 2023). We select the method proposed by (Zhong et al., 2017) in our theoretical analysis since it can provably achieve a relatively small error with least assumptions.
Assumption A.1.
@@
  • Data: Let data matrix $X:=[{\bm{x}}_{1},\cdots,{\bm{x}}_{B}]\in\mathbb{R}^{d\times B}$, we denote the $B$-th singular value by $\pi_{\min}>0$. Training samples are normalized: $\|{\bm{x}}_{i}\|=1,\forall i\in[B]$.
@@
We first introduce the method in (Zhong et al., 2017) when the activation function satisfies $k_{2}=2$ and $k_{3}=3.$ In addition to $\hat{{\bm{T}}}=\sum_{j=1}^{m}g({\bm{w}}_{j})H_{3}({\bm{w}}_{j})$, we also need the matrix $\hat{P}=\sum_{j=1}^{m}g({\bm{w}}_{j})H_{2}({\bm{w}}_{j})$ in the tensor decomposition, where $H_{p}$ is the $p$-th Hermite function. We first conduct the power method to $\hat{P}$ to estimate the orthogonal span $U$ of training samples ${\bm{x}}_{1},\dots,{\bm{x}}_{B}$ and denote it by $V$, where $V\in\mathbb{R}^{B\times d}$ is an orthogonal matrix. Then we conduct tensor decomposition with $\hat{{\bm{T}}}(V,V,V)$ instead of $T$ and have the estimation of $\{V^{\top}{\bm{x}}_{i}\}_{i=1}^{B}$. By multiplying column orthogonal matrix $V$, we can reconstruct training data. The advantage of this method is that the dimension of $\hat{{\bm{T}}}(V,V,V)$ is $B<d$. Then the error will depend on $B$ instead of $d$.
@@ A.2 Error Bound
Theorem A.1 (Adapted from the proof of Theorem 5.6 in (Zhong et al., 2017).).
@@
where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ satisfying Assumption A.1. Let $V$ be the output of Algorithm 3 in (Zhong et al., 2017) with input $P$ and $\{s_{i}{\bm{u}}_{i}\}_{i=1}^{B}$ be the output of Algorithm 1 in (Kuleshov et al., 2015) with input ${\bm{T}}(V,V,V)$, where $\{s_{i}\}$ are unknown signs. Suppose the perturbations satisfy
@@
Let $N=\Theta(\log\frac{1}{\epsilon})$ be the iteration numbers of Algorithm 3 in (Zhong et al., 2017), where $\epsilon=\frac{\mu}{\nu_{\min}}$. Then with high probability, we have
@@
With Theorem A.1, we only need to bound the error $\|P-\hat{P}\|$ and $\|{\bm{T}}-\hat{{\bm{T}}}\|$, where the bound varies under different settings. Wang et al. (2023) proved the error bound under the square loss setting. In the following section, we will propose error bounds with cross-entropy loss, which is more widely used in classification problems. We will also give error bounds when defending strategies are used against privacy attacks.
A.3 Matrix Bernstein’s Inequality
@@
Lemma A.2 (Matrix Bernstein for unbounded matrices; adapted from Lemma B.7 in (Zhong et al., 2017)).
@@
Appendix B Analysis of Feature Reconstruction
Wang et al. (2023) proposed a privacy attack based on tensor decomposition but they only analyzed the case of square loss. However, in classification problems, cross-entropy loss is widely used. Here we propose a conduct feature reconstruction method with tensor decomposition under cross-entropy loss setting, which is the detailed version of the method proposed in Section 2.1.
We define the network the same as Section 2.1. Let the input dimension be $d$ and the number of classes is $K$. For a two layer neural network $f({\bm{x}};\Theta)=\frac{1}{m}A\sigma(W{\bm{x}})$ with $m$ hidden nodes, where $W=[{\bm{w}}_{1},\dots,{\bm{w}}_{m}]\in\mathbb{R}^{m\times d}$, $A=[{\bm{a}}_{1},\dots,{\bm{a}}_{K}]^{\top}\in\mathbb{R}^{K\times m}$ and $\sigma$ is point-wise. Denote $z_{k}=a_{k}^{\top}\sigma(W{\bm{x}})$ and $p_{k}=\mathrm{Softmax}(z_{k})$. Then cross entropy loss is $\ell(f({\bm{x}};\Theta),{\bm{y}})=-\sum_{k=1}^{K}y_{k}\log(p_{k})$, where ${\bm{y}}\in\mathbb{R}^{K}$ is a one-hot vector. Then we have the gradient
$\frac{\partial\ell}{\partial{\bm{w}}_{j}}=\sum_{k=1}^{K}\frac{\partial\ell}{\partial z_{k}}\frac{\partial z_{k}}{\partial{\bm{w}}_{j}}=\frac{1}{m}\sum_{k=1}^{K}(p_{k}-y_{k})a_{kj}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}){\bm{x}}$
@@ Appendix B Analysis of Feature Reconstruction
$\frac{\partial\ell}{\partial a_{kj}}=\frac{\partial\ell}{\partial z_{k}}\frac{\partial z_{k}}{\partial a_{kj}}=\frac{1}{m}(p_{k}-y_{k})\sigma({\bm{w}}_{j}^{\top}{\bm{x}}).$
To reconstruct training data, we can modify the initialization weights arbitrarily. Let ${\bm{a}}_{k}=c_{k}\tilde{{\bm{a}}}$, where $c_{k}$ is a scalar and $\tilde{{\bm{a}}}\sim\mathcal{N}(0,I_{m})$. Then each rows of $A$ are proportional so we can write $z_{k}=c_{k}z$. We can let $c_{k}=b_{1}$ for $k\in U$, where $U$ is a subset of $\{1,\dots,K\}$ and $\|U\|=K/2$ (assuming $K>2$ is even) and $c_{k}=b_{2}$ for $k\notin{U}$, where $b_{1}\neq b_{2}$. Then all $p_{k}$’s are equal to some $\tilde{p}$ for $k\in U$ and all $p_{k}$’s are equal to another $\tilde{q}$ for $k\notin U$. Without loss of generality, we can assume that $1\in U$ and $2\notin U$, then $c_{1}=b_{1}$ and $c_{2}=b_{2}$. Note that
$p_{1}+p_{2}=\tilde{p}+\tilde{q}=\frac{e^{b_{1}z}+e^{b_{2}z}}{\sum_{k\in U}e^{b_{1}z}+\sum_{k\notin U}e^{b_{2}z}}=\frac{2}{K}$
is a constant, no matter the input samples.
If we use the gradient of $A$, we construct $g({\bm{w}}_{j})$ with the first two classes:
$\begin{split}g({\bm{w}}_{j}):&=\sum_{i=1}^{B}\left(\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{1j}}+\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{2j}}\right)\\ &=\frac{1}{m}\sum_{i=1}^{B}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})(p_{1}-y_{i1}+p_{2}-y_{i2})\\ &=\frac{1}{m}\sum_{i=1}^{B}\left(\frac{2}{K}-(y_{i1}+y_{i2})\right)\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i}).\end{split}$ (8)

### v2

@@ 6 Conclusion
In this paper, we systematically evaluate various defense methods, both theoretically and empirically. We established the algorithmic upper bound and matching information-theoretic lower bound of the reconstruction error for various defenses on two-layer random nets, thereby quantitatively analyzing the effect of defense strength, data dimension, and model width. To extend our theoretical insights to more general network architectures and explore the utility-privacy trade-off, we introduced a robust attack method capable of overcoming a broad spectrum of defense strategies in a setting with an honest but curious server. The method enhances the conventional gradient inversion attack by combining it with the powerful feature reconstruction attack that achieves our algorithmic upper bound. We propose to evaluate each attack based on their impact on utility and effectiveness against the strongest corresponding attack. Our evaluation indicates gradient pruning is the strongest among our considered defenses and against our considered attacks, which updates some previous evaluation results that are purely based on gradient inversion attacks (Huang et al., 2021). Our work establishes a more comprehensive understanding of the data reconstruction problem in federated learning. We anticipate that our proposed framework will encourage further research and additional evaluations building on our findings.
Acknowledgement
This work was supported by the U.S. Department of Energy, Office of Science, under Award Number DE-SC0024721 and the Whitehead Fellowship from New York University. QL would like to thank Yanjun Han for the valuable discussions on the lower bound formulation of the paper.
Appendix A Analysis of Differential Privacy
Differential privacy (DP) [Dwork, 2006] is a measure of privacy mainly in membership inference attack. A random algorithm $\mathcal{M}:\mathcal{D}\rightarrow\mathcal{R}$ satisfies $(\epsilon,\delta)$-DP if for any $\mathcal{Q}\subset\mathcal{R}$ and any adjacent inputs $S,S^{\prime}\in\mathcal{D}$ holds that
$\mathbb{P}(\mathcal{M}(S)\in\mathcal{Q})\leq e^{\epsilon}\mathbb{P}(\mathcal{M}(S^{\prime})\in\mathcal{Q})+\delta.$
Here adjacent inputs means only one of the samples is different.
In reconstruction attack, DP is too strong that an algorithm may prevent data leakage without DP guarantees. Moreover, the price to guarantee DP is high. A popular method to guarantee DP is DP-SGD [Abadi et al., 2016]. However, it requires extremely large Gaussian noise.
Proposition A.1 (Full version of Proposition 2.1).
We define a two-layer neural network $f({\bm{x}};\Theta)=\sum_{j=1}^{m}a_{j}\sigma({\bm{w}}_{j}^{\top}{\bm{x}})$ with input dimension $d$, random Gaussian weights $a_{j}\sim\mathcal{N}(0,\frac{1}{m^{2}})$ and ${\bm{w}}_{j}\sim\mathcal{N}(0,I_{d})$ and activation function $\sigma$ is 1-Lipschitz. The loss function $\ell$ is square loss, the data $\|{\bm{x}}\|=1$ and the label $y\in\{\pm 1\}$. We denote the gradient of $\ell$ by $G$. The randomized mechanism $\mathcal{M}=G+\mathcal{N}(0,\sigma^{2}I)$ is $(\epsilon,\delta)-$ DP for any $\epsilon,\delta>0$ if $\sigma^{2}=\Omega(\frac{m\log(1/\delta)}{\epsilon})$ with high probability with respect to random weights.
In order to prove the proposition, we first define
$\alpha_{\mathcal{M}}(\lambda):=\max_{S,S^{\prime}}\log\operatorname{\mathbb{E}}_{z\sim\mathcal{M}(S)}\left(\frac{p(\mathcal{M}(S)=z)}{p(\mathcal{M}(S^{\prime})=z)}\right)^{\lambda},$
where $p$ is the density function. In DP-SGD setting, $\mathcal{M}(S)=G(S)+\mathcal{M}(0,\sigma^{2}I)$ so $\mathcal{M}(S)\sim\mathcal{N}(G(S),\sigma^{2}I)$ for $S\in\mathbb{R}^{d\times B}$. The following lemma is a key step in the proof of Proposition A.1.
Lemma A.2 (Theorem 2 in [Abadi et al., 2016]).
For any $\epsilon>0$, the mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP for $\delta=\min_{\lambda}(\alpha_{\mathcal{M}}(\lambda)-\lambda\epsilon)$.
Proof of Proposition A.1.
We denote $\Delta:=\max_{{\bm{x}},{\bm{x}}^{\prime}\in\mathbb{R}^{d}}\|G({\bm{x}})-G)({\bm{x}}^{\prime})\|^{2}$. Then we compute $\alpha_{\mathcal{M}}(\lambda)$ with adjacent $S,S^{\prime}\in\mathbb{R}^{d\times B}$. We denote the only different element by ${\bm{x}}$ and ${\bm{x}}^{\prime}$.
$\alpha_{\mathcal{M}}(\lambda)=$ $\max_{S,S^{\prime}}\log\operatorname{\mathbb{E}}_{Z\sim\mathcal{N}(G(S),\sigma^{2}I)}\left[\left(\frac{p_{S}(Z)}{p_{Y}(Z)}\right)^{\lambda}\right]$
$=$ $\max_{{\bm{x}},{\bm{x}}^{\prime}}\log\operatorname{\mathbb{E}}_{{\bm{z}}\sim\mathcal{N}(G({\bm{x}}),I)}\exp\left(\frac{\lambda}{2\sigma^{2}}\sum_{i=1}^{d}\left(\left(G({\bm{x}})_{i}-z_{i}\right)^{2}-\left(G({\bm{x}}^{\prime})_{i}-z_{i}\right)^{2}\right)\right)$
$=$ $\max_{{\bm{x}},{\bm{x}}^{\prime}}\log\frac{1}{\left(\sqrt{2\pi}\sigma\right)^{d}}\int\exp\left(\frac{1}{2\sigma^{2}}\sum_{i=1}^{d}\left((\lambda+1)\left(G({\bm{x}})_{i}-z_{i}\right)^{2}-\lambda\left(G({\bm{x}}^{\prime})_{i}-z_{i}\right)^{2}\right)\right)$
$=$ $\max_{{\bm{x}},{\bm{x}}^{\prime}}\log\frac{1}{\left(\sqrt{2\pi}\sigma\right)^{d}}\int\exp\left(\frac{1}{2\sigma^{2}}\sum_{i=1}^{d}\left(\left(z_{i}-\left((\lambda+1)G({\bm{x}})_{i}-\lambda G({\bm{x}}^{\prime})_{i}\right)\right)^{2}-\lambda(\lambda+1)\left(G({\bm{x}})_{i}-G({\bm{x}}^{\prime})_{i}\right)^{2}\right)\right)$
$=$ $\max_{{\bm{x}},{\bm{x}}^{\prime}}\log\exp\left(\frac{\lambda(\lambda+1)}{2\sigma^{2}}\sum_{i=1}^{d}\left(G({\bm{x}})_{i}-G({\bm{x}}^{\prime})_{i}\right)^{2}\right)$
$=$ $\max_{{\bm{x}},{\bm{x}}^{\prime}}\frac{\lambda(\lambda+1)}{2\sigma^{2}}\left\|G({\bm{x}})-G({\bm{x}}^{\prime})\right\|^{2}=\frac{\lambda(\lambda+1)}{w\sigma^{2}}\Delta.$
By Lemma A.2, for any
$\delta\geq$ $\min_{\lambda}\exp\left(\frac{\lambda(\lambda+1)}{2\sigma^{2}}\Delta-\lambda\epsilon\right)$
$=\exp\left(-\frac{\Delta}{2\sigma^{2}}\left(\frac{\sigma^{2}\epsilon}{\Delta}-\frac{1}{2}\right)^{2}\right),$
$\mathcal{M}$ is $(\epsilon,\delta)$-DP. Then for $\sigma^{2}\geq\Omega(\frac{\Delta\log(1/\delta)}{\epsilon})$, $\mathcal{M}$ is $(\epsilon,\delta)$-DP.
Now we only need to bound $\Delta$. We denote $r=y-f({\bm{x}})$ and $r^{\prime}=y^{\prime}-f({\bm{x}}^{\prime})$. We consider $\nabla_{a_{j}}$ and $\nabla_{{\bm{w}}_{j}}$ separately. For $\nabla_{a_{j}}$, we have
$\left|\nabla_{a_{j}}\ell(S)-\nabla_{a_{j}}\ell(S^{\prime})\right|^{2}=$ $\left|\nabla_{a_{j}}\ell({\bm{x}})-\nabla_{a_{j}}\ell({\bm{x}}^{\prime})\right|^{2}$
$=$ $\left|r\sigma({\bm{w}}_{j}^{\top}{\bm{x}})-r^{\prime}\sigma({\bm{w}}_{j}^{\top}{\bm{x}})\right|^{2}$
$\lesssim$ $r^{2}\left|\sigma({\bm{w}}_{j}^{\top}{\bm{x}})\right|^{2}+r^{\prime 2}\left|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}^{\prime})\right|^{2}$
$=$ $\tilde{O}(1)$
with high probability for any $j$. For $\nabla_{{\bm{w}}_{j}}$, we have
$\left\|\nabla_{{\bm{w}}_{j}}\ell(S)-\nabla_{{\bm{w}}_{j}}\ell(S^{\prime})\right\|^{2}=$ $\left\|\nabla_{{\bm{w}}_{j}}\ell({\bm{x}})-\nabla_{{\bm{w}}_{j}}\ell({\bm{x}}^{\prime})\right\|^{2}$
$=$ $\left\|r\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}){\bm{x}}-r^{\prime}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}^{\prime}){\bm{x}}^{\prime}\right\|^{2}$
$\lesssim$ $r^{2}\left\|{\bm{x}}\right\|^{2}+r^{\prime 2}\left\|{\bm{x}}^{\prime}\right\|^{2}$
$=$ (~$\tilde{(}O)(1)$
with high probability for any $j$. Then
$\Delta$ $=\max_{S,S^{\prime}}\left\|G(S)-G(S^{\prime})\right\|^{2}$
$=\max_{S,S^{\prime}}\sum_{j=1}^{m}\left(\left|\nabla_{a_{j}}\ell(S)-\nabla_{a_{j}}\ell(S^{\prime})\right|^{2}+\left\|\nabla_{{\bm{w}}_{j}}\ell(S)-\nabla_{{\bm{w}}_{j}}\ell(S^{\prime})\right\|^{2}\right)$
$\leq\tilde{O}(m)$
with high probability. Therefore, for $\sigma^{2}\geq\tilde{\Omega}(\frac{m\log(1/\delta)}{\epsilon})$, $\mathcal{M}$ is $(\epsilon,\delta)$-DP with high probability. ∎
Appendix B Analysis of Reconstruction Upper Bound
B.1 Tensor Method
@@
Various of noisy tensor decomposition can be used in tensor based attack [Wang et al., 2023]. We select the method proposed by [Zhong et al., 2017] in our theoretical analysis since it can provably achieve a relatively small error with least assumptions.
Assumption B.1.
@@
  • Data: Let data matrix $S:=[{\bm{x}}_{1},\cdots,{\bm{x}}_{B}]\in\mathbb{R}^{d\times B}$, we denote the $B$-th singular value by $\pi_{\min}>0$. Training samples are normalized: $\|{\bm{x}}_{i}\|=1,\forall i\in[B]$.
@@
We first introduce the method in [Zhong et al., 2017] when the activation function satisfies $k_{2}=2$ and $k_{3}=3.$ In addition to $\hat{{\bm{T}}}=\sum_{j=1}^{m}g({\bm{w}}_{j})H_{3}({\bm{w}}_{j})$, we also need the matrix $\hat{P}=\sum_{j=1}^{m}g({\bm{w}}_{j})H_{2}({\bm{w}}_{j})$ in the tensor decomposition, where $H_{p}$ is the $p$-th Hermite function. We first conduct the power method to $\hat{P}$ to estimate the orthogonal span $U$ of training samples ${\bm{x}}_{1},\dots,{\bm{x}}_{B}$ and denote it by $V$, where $V\in\mathbb{R}^{B\times d}$ is an orthogonal matrix. Then we conduct tensor decomposition with $\hat{{\bm{T}}}(V,V,V)$ instead of $T$ and have the estimation of $\{V^{\top}{\bm{x}}_{i}\}_{i=1}^{B}$. By multiplying column orthogonal matrix $V$, we can reconstruct training data. The advantage of this method is that the dimension of $\hat{{\bm{T}}}(V,V,V)$ is $B<d$. Then the error will depend on $B$ instead of $d$.
@@ B.2 Error Bound
Theorem B.1 (Adapted from the proof of Theorem 5.6 in [Zhong et al., 2017].).
@@
where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ satisfying Assumption B.1. Let $V$ be the output of Algorithm 3 in [Zhong et al., 2017] with input $P$ and $\{s_{i}{\bm{u}}_{i}\}_{i=1}^{B}$ be the output of Algorithm 1 in [Kuleshov et al., 2015] with input ${\bm{T}}(V,V,V)$, where $\{s_{i}\}$ are unknown signs. Suppose the perturbations satisfy
@@
Let $N=\Theta(\log\frac{1}{\epsilon})$ be the iteration numbers of Algorithm 3 in [Zhong et al., 2017], where $\epsilon=\frac{\mu}{\nu_{\min}}$. Then with high probability, we have
@@
With Theorem B.1, we only need to bound the error $\|P-\hat{P}\|$ and $\|{\bm{T}}-\hat{{\bm{T}}}\|$, where the bound varies under different settings. Wang et al., 2023 proved the error bound under the square loss setting. In the following section, we will propose error bounds with cross-entropy loss, which is more widely used in classification problems. We will also give error bounds when defending strategies are used against privacy attacks.
B.3 Matrix Bernstein’s Inequality
@@
Lemma B.2 (Matrix Bernstein for unbounded matrices; adapted from Lemma B.7 in [Zhong et al., 2017]).
@@
B.4 Reconstruction Upper Bounds
Wang et al., 2023 proposed a privacy attack based on tensor decomposition. We adopt the setting of that work. Let the input dimension be $d$ For a two layer neural network $f({\bm{x}};\Theta)=\sum_{j=1}^{m}a_{j}\sigma({\bm{w}}_{j}^{\top}{\bm{x}})$ with $m$ hidden nodes and $\sigma$ is point-wise. Then square loss with $B$ samples is $\ell(f({\bm{x}};\Theta),y)=\sum_{i=1}^{B}(y_{i}-f({\bm{x}}_{i};\Theta))^{2}$. Then we have the gradient
$\frac{\partial\ell}{\partial{\bm{w}}_{j}}=\sum_{i=1}^{B}r_{i}a_{j}\sigma^{\prime}({\bm{w}}_{j}\top{\bm{x}}_{i}){\bm{x}}_{i}$
@@ B.4 Reconstruction Upper Bounds
$\frac{\partial\ell}{\partial a_{j}}=\sum_{i=1}^{B}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i}),$
where $r_{i}=2(f({\bm{x}}_{i};\Theta)-y_{i})$.
To reconstruct training data, we set the weights to be random Gaussian.
Assumption B.2.
The network $f({\bm{x}};\Theta)=\sum_{j=1}^{m}a_{j}\sigma({\bm{w}}_{j}^{\top}{\bm{x}})$. The parameters of the network satisfy $a_{j}\sim\mathcal{N}(0,\frac{1}{m^{2}})$ and ${\bm{w}}_{j}\sim\mathcal{N}(0,I_{d})$ and are independent.